The area of a rectangle is represented by n^2-n-12. If the length of the rectangle is represented by n+3, express the width of the rectangle in terms of n.
Found 3 solutions by unlockmath, mananth, ewatrrr:
Answer by unlockmath(1688)

Hello,
We know that area is calculated by length times height, right. So we can write this equation as:
n^2-n-12
This can be factored as:
(n-4)(n+3)
Width can be (n-4)
RJ
www.math-unlock.com

Answer by mananth(16949)

The area of a rectangle is represented by n^2-n-12. If the length of the rectangle is represented by n+3, express the width of the rectangle in terms of n.
Area = n2^-n-12
length = n+3
..
width = n^2-n-12/(n+3)
width = (n-4)(n+3)/(n+3)
width = (n-4)
...
m.ananth@hotmail.ca

Answer by ewatrrr(24785)


Hi,
Area of a rectangle = Length*Width
question states***
A = n^2-n-12 = L*W Question states L = (n+3)
factoring
(n+3)(n-4)= L*W therefore W can be expressed as (n-4)